# The rationality problem for multinorm one tori

### v2

@@ 1. Introduction
We remark that the direction of the implication cannot be reversed. See Section 2 for detail.
@@
The rationality problem for norm one tori has been extensively investigated by [EM75], [CS77], [Hür84], [CS87], [LeB95], [CK00], [LL00], [Flo], [End11], [HY17], [HHY20], [HY21], [HY24] and [HY]. On the other hand, the rationality problem for multinorm one tori (especially not norm one tori) has been studied by [Hür84], [End01]. Moreover it has been treated in [CS77], [CK00], [End11]. However, less is known about it compared to the norm one tori.
@@ 1.1. Main theorems
      $r^{\prime}=1$ and $L^{\prime}/k$ which is cyclic over $k$; or
@@ 1.2. Application to multinorm principle
Theorem 1.9 gives a new example on the validity of the multinorm principle. It follows from Theorem 1.5 together with the standard argument on the multinorm principle. See Section 9 for more details.
@@
In Section 2, we prepare some basic definitions and known results about the rationality of algebraic tori. In particular, we discuss the relationship between algebraic tori and $G$-lattices. In Section 3, we introduce the concept of multinorm tori, and provide a generalization of their corresponding $G$-lattices. Furthermore, we develop the technique of Endo [End11], and construct some reduction methods for an investigation of the rationality problem for arbitrary multinorm one tori. In Section 4, we review some properties of $p$-groups used in this paper. In Section 5, we give a proof of Theorem 1.2. In Section 6, for certain $G$-lattices, we determine whether they are stably permutation or not quasi-invertible using the theory of flabby resolutions. These lattices play a crucial role in Section 7 and Section 8. In Section 7, we give a proof of Theorem 1.3 by dividing into four steps. In Section 8, we give a proof of Theorem 1.5. That is, we give a necessary and sufficient condition for the multinorm one tori to be stably rational and retract rational in the case that split over finite Galois extensions with nilpotent Galois groups. Finally, Section 9 gives a proof of Theorem 1.9, which gives an application of our study on the rationality problem for multinorm one tori.
@@ 2. Basic facts on the rationality of tori
In this paper, we concentrate on the case where $X$ is an algebraic torus over $k$. Fix a separable closure $k^{\operatorname{sep}}$ of $G$. Then, we can rephrase the stable rationality and the retract rationality of algebraic tori by means of $G$-lattices, where $G$ is a finite quotient of the Galois group of $k^{\operatorname{sep}}/k$. We follow the same terminology as the text book [Lor05] and the mainly referenced paper [End11].
@@
    where $P$ is permutation and $F$ is coflabby.
@@
    where $P$ is permutation and $F$ is flabby.
@@
By assumption, $F$ is stably permutation. Hence, there is a permutation $G$-lattice $R$ such that $F\oplus R^{\prime}$ is permutation. Moreover, the sequence
@@
Let $G$ be a finite group, and $H$ its subgroup. Consider a $G$-lattice $M$. If $M$ is a quasi-permutation (resp. quasi-invertible) $G$-module, then so is as an $H$-lattice.
@@
    If the $G$-lattice $M$ is quasi-permutation (resp. quasi-invertible), then so is for the $G/N$-lattice $M^{[N]}$.
@@
For a $G$-lattice $M$, we can be summarized as follows:
@@ 3.1. Multinorm one tori
In this paper, we determine whether the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation or quasi-invertible in order to classify the stably/retract rationality of multinorm one tori. In the sequel of this section, we discuss
@@ 3.3. Reduction to smaller $G$-lattices
where $m\in\mathbb{Z}_{>0}$ and $c_{1},\ldots,c_{m}$ are integers with great common divisor $d$. Then there is an automorphism $\lambda$ of $G$-lattices $\mathbb{Z}[G/H]^{\oplus m}$ such that $f\circ\lambda$ coincides with the composite
@@ 4. $p$-groups
In the sequel of this paper, for a positive integer $n$, we denote by $D_{n}$ the dihedral group of order $2n$, that is,
@@
(i) $\Rightarrow$ (ii) is clear. (iii) $\Rightarrow$ (i) follows from Proposition 5.2. In the following, we prove (ii) $\Rightarrow$ (iii), which is archived by giving a proof of the contraposition. We may assume that $\mathcal{H}$ contains at least two elements. In particular, $G$ is not cyclic according to Lemma 5.7. It suffices to prove that $J_{G/H}$ is not quasi-invertible if $\mathcal{H}\geq 2$. Write $G=p^{\nu}$. We give a proof of the above assertion by induction on $\nu$. If $\nu=2$, the assertion follows from Proposition 5.3. Now suppose $\nu\geq 3$, and the assertion holds for all $\nu-1$. If $N^{G}(\mathcal{H})\neq\{1\}$, take a subgroup $N$ of order $p$ in $Z(G)\cap N^{G}(\mathcal{H})$. Note that $Z(G)\cap N^{G}(\mathcal{H})$ is non-trivial by Lemma 4.3. Then it suffices to prove the assertion for the $G/N$-lattice $J_{(G/N)/\mathcal{H}^{N}}$. Since $(G/N)=p^{\nu-1}$, the assertion follows from the induction hypothesis. Hence, we may further assume $N^{G}(\mathcal{H})=\{1\}$. If $M(\mathcal{H})=p$, then $G$ is elementary $p$-abelian since $N^{G}(\mathcal{H})=\{1\}$. Therefore, the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible by Proposition 5.3. Therefore, we can impose $M(\mathcal{H})\geq p^{2}$ in the sequel. By the induction hypothesis, it suffices to prove that there is a maximal subgroup $P$ of $G$ that satisfies $\mathcal{H}_{P}^{\operatorname{red}}\geq 2$.
@@ 6.1. Quasi-permutation lattices
where the horizontal sequences are exact. Note that the central and the rightmost vertical maps are injective, and the cokernel of the central vertical map is torsion-free. Moreover, since the the cokernel of $\Phi_{m,\nu,\ast}$ is torsion, we have
@@ 7.3. Third step: The case admitting normal factors
Case 2-c. $\Sigma=\sigma$. Let $P$ be a maximal subgroup of $G$ containing $N_{G}(H_{1})$. It suffices to prove that $\mathcal{H}_{P}^{\operatorname{red}}$ is not a subset of $\{gH_{1}g^{-1}\mid g\in G\}$. This implies $\mathcal{H}_{P}^{\operatorname{red}}\geq 3$ since $H_{1}$ is not normal in $G$. Assume not, that is, $\mathcal{H}_{P}^{\operatorname{red}}$ is contained in $\{gH_{1}g^{-1}\mid g\in G\}$. Since $H_{0}$ is normal of index $\Sigma=\sigma=\mu(\mathcal{H})$, it is contained in $N_{G}(H_{1})$ by Lemma 5.9. In particular, $H_{0}$ is an element of $\mathcal{H}_{P}^{\operatorname{red}}$, which contradicts to the assumption. Hence the proof is complete. $\blacksquare$
@@ 7.4. Final step: General case
      $G/N^{G}(\mathcal{H})\cong D_{2^{\nu}}$ and $\mathcal{H}=\{\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle,\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle\}$ for some $\nu\in\mathbb{Z}_{>0}$ and $m,m^{\prime}\in\mathbb{Z}$.
@@
(ii) $\Rightarrow$ (iii): We may assume that $\mathcal{H}$ is strongly reduced. For each odd prime divisor $p$ of #G\#G, we denote by $G_{p}$ the unique $p$-Sylow subgroup of $G$. Then $J_{G/\mathcal{H}}$ is quasi-invertible as a $G_{p}$-lattice. On the other hand, the set $\mathcal{H}_{G_{p}}^{\operatorname{set}}$ consists of $G_{p}\cap H$ for all $H\in\mathcal{H}$. Hence, by Theorem 5.10, there is $H_{p}\in\mathcal{H}$ so that $N_{p}:=G_{p}\cap H_{p}\triangleleft G_{p}$, $G_{p}/N_{p}$ is cyclic and $G_{p}\cap H\subset N_{p}$ for any $H\in\mathcal{H}$. In particular, the subgroup $N_{p}$ is normal in $G$. Because $G$ is nilpotent, it is a product of $G_{p}$ for all prime divisors $p$ of #G\#G, Lemma 8.3 implies the existence of a strongly reduced set of subgroups $\mathcal{H}^{\prime}$ of $G$ so that
  • $(\mathcal{H}^{\prime})_{G_{p}}^{\operatorname{set}}=\{N_{p}\}$ for any odd prime $p$, where $N_{p}\triangleleft G_{p}$ and $G_{p}/N_{p}$ is cyclic; and
  • $J_{G/\mathcal{H}}]=[J_{G/\mathcal{H}^{\dagger}}]$.
Therefore, by replacing $\mathcal{H}$ and $G$ with $\mathcal{H}^{\dagger}$ and $G/N^{G}(\mathcal{H}^{\dagger})$, respectively, we may assume $G=C_{m}\times G_{2}$ for some odd integer $m$ and #H\#H is a power of $2$ for every $H\in\mathcal{H}$. In this case, the set of subgroups $\mathcal{H}_{G_{2}}^{\operatorname{set}}$ of $G_{2}$ is strongly reduced since $\mathcal{H}$ is so. Hence, the assumption that $J_{G/\mathcal{H}^{\prime\prime}}$ is quasi-invertible implies that
  1. (A)
    $\mathcal{H}_{G_{2}}^{\operatorname{set}}=\{N_{2}\}$ for some $N\triangleleft G_{2}$ so that $G_{2}/N_{2}$ is cyclic; or
  2. (B)
    $\mathcal{H}_{G_{2}}^{\operatorname{set}}=\{H,H^{\prime}\}$, $G_{2}/N^{G_{2}}(\mathcal{H}_{G_{2}}^{\operatorname{set}})\cong D_{2^{\nu}}$ for some $n\in\mathbb{Z}_{>0}$, $H/N^{G_{2}}(\mathcal{H}_{G_{2}}^{\operatorname{set}})\cong\langle\tau_{2^{\nu}}\rangle$ and $H^{\prime}/N^{G_{2}}(\mathcal{H}_{G_{2}}^{\operatorname{set}})\cong\langle\sigma_{2^{\nu}}\tau_{2^{\nu}}\rangle$.
If (A) holds, then the set $\mathcal{H}^{\prime}:=\{N_{2}\}$ satisfies (iii-$\alpha$). On the other hand, if (B) is valid, then the set $\mathcal{H}^{\prime}:=\{H,H^{\prime}\}$ is the one that satisfies (iii-$\beta$).
@@
(ii) $\Rightarrow$ (iii): Combining the assumption with Proposition 3.1, we obtain that the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-invertible. Hence, there exists a reduced set of subgroups in $G$ with $\mathcal{H}^{\prime}<\mathcal{H}$ and $J_{G/\mathcal{H}^{\prime}}]=[J_{G/\mathcal{H}}]$ that satisfies (iii-$\alpha$) or (iii-$\beta$) in Theorem 8.4. If (iii-$\alpha$) holds, denote by $\mathbf{K}^{\prime}=K^{\prime}$ the intermediate field of $\widetilde{K}/k$ corresponding to $N$. Then, $K^{\prime}$ is a subalgebra of $\mathbf{K}$ satisfying (iii-a) in Theorem 1.5. On the other hand, if (iii-$\beta$) is true, write $\mathcal{H}=\{H,H^{\prime}\}$, and denote by $K^{\prime}_{1}$ and $K^{\prime}_{2}$ the intermediate fields of $L/k$ corresponding to $K^{\prime}_{1}$ and $K^{\prime}_{2}$, respectively. Let $\mathbf{K}^{\prime}:=K^{\prime}_{1}\times K^{\prime}_{2}$, then it is a subalgebra of $\mathbf{K}$ satisfying (iii-b) in Theorem 1.5.
@@
$^{1}(k,T):=\operatorname{Ker}\left(H^{1}(k,T)\xrightarrow{(\operatorname{Res}_{k_{v}/k})_{v}}\bigoplus_{v}H^{1}(k_{v},T)\right).$
@@
Here, $\operatorname{Pic}(\overline{X})^{\vee}$ and $A_{k}(T)$ is defined as follows:

### v3

@@ 1. Introduction
We remark that the direction of the implication cannot be reversed. See Section 2 for more details.
@@
The rationality problem for norm one tori has been extensively investigated by [EM75], [CS77], [Hür84], [CS87], [LeB95], [CK00], [LL00], [Flo], [End11], [HY17], [HHY20], [HY21], [HY24] and [HY]. However, less is known about the rationality problem for multinorm one tori compared to the norm one tori. It has been studied by [Hür84], [End01]. Moreover it has been treated in [CS77], [CK00], [End11].
@@ 1.1. Main theorems
      $r^{\prime}=1$ and $L^{\prime}/k$ is cyclic over $k$; or
@@ 1.2. Application to multinorm principle
Theorem 1.9 gives a new example of the validity of the multinorm principle. It follows from Theorem 1.5 together with the standard argument on the multinorm principle. Note that our proof uses an extension of Ono’s theorem ([Ono63]), that is, a connection between $\Sha(\mathbf{K}/k)$ and the Tate–Shafarevich group of $T_{\mathbf{K}/k}$. See Section 9 for more details.
@@
In Section 2, we prepare some basic definitions and known results about the rationality of algebraic tori. In particular, we discuss the relationship between algebraic tori and $G$-lattices. In Section 3, we introduce the concept of multinorm tori, and provide a generalization of their corresponding $G$-lattices. Furthermore, we develop the technique of Endo [End11], and construct some reduction methods to investigate the rationality problem for arbitrary multinorm one tori. In Section 4, we review some properties of $p$-groups used in this paper. In Section 5, we give a proof of Theorem 1.2. In Section 6, for certain $G$-lattices, we determine whether they are stably permutation or not quasi-invertible using the theory of flabby resolutions. These lattices play a crucial role in Section 7 and Section 8. In Section 7, we give a proof of Theorem 1.3 by dividing into four steps. In Section 8, we give a proof of Theorem 1.5. That is, we give a necessary and sufficient condition for the multinorm one tori to be stably rational and retract rational in the case that split over finite Galois extensions with nilpotent Galois groups. Finally, Section 9 gives a proof of Theorem 1.9, which gives an application of our study on the rationality problem for multinorm one tori.
@@ 2. Basic facts on the rationality of tori
In this paper, we concentrate on the case where $X$ is an algebraic torus over $k$. Fix a separable closure $k^{\operatorname{sep}}$ of $k$. Then, we can rephrase the stable rationality and the retract rationality of algebraic tori by means of $G$-lattices, where $G$ is a finite quotient of the Galois group of $k^{\operatorname{sep}}/k$. We follow the same terminology as the text book [Lor05] and the mainly referenced paper [End11].
@@
    where $R$ is permutation and $U$ is coflabby.
@@
    where $R$ is permutation and $F$ is flabby.
@@
By assumption, $F$ is stably permutation. Hence, there is a permutation $G$-lattice $R^{\prime}$ such that $F\oplus R^{\prime}$ is permutation. Moreover, the sequence
@@
Let $G$ be a finite group, and $H$ its subgroup. Consider a $G$-lattice $M$. If $M$ is a quasi-permutation (resp. quasi-invertible) $G$-module, then $M$ is so as an $H$-lattice.
@@
    If the $G$-lattice $M$ is quasi-permutation (resp. quasi-invertible), then the $G/N$-lattice $M^{[N]}$ is so.
@@
For a $G$-lattice $M$, we can summarize the properties mentioned above as follows:
@@ 3.1. Multinorm one tori
In this paper, we determine whether the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-permutation or quasi-invertible in order to classify the stably/retract rationality of multinorm one tori. In what follows, we discuss
@@ 3.3. Reduction to smaller $G$-lattices
where $m\in\mathbb{Z}_{>0}$ and $c_{1},\ldots,c_{m}$ are integers with the greatest common divisor $d$. Then there is an automorphism $\lambda$ of $G$-lattices $\mathbb{Z}[G/H]^{\oplus m}$ such that $f\circ\lambda$ coincides with the composite
@@ 4. $p$-groups
In what follows, for a positive integer $n$, we denote by $D_{n}$ the dihedral group of order $2n$, that is,
@@
(i) $\Rightarrow$ (ii) is clear. (iii) $\Rightarrow$ (i) follows from Proposition 5.2. In the following, we prove (ii) $\Rightarrow$ (iii), which is achieved by giving a proof of the contraposition. We may assume that $\mathcal{H}$ contains at least two elements. In particular, $G$ is not cyclic according to Lemma 5.7. It suffices to prove that $J_{G/H}$ is not quasi-invertible if $\mathcal{H}\geq 2$. Write $G=p^{\nu}$. We give a proof of the above assertion by induction on $\nu$. If $\nu=2$, the assertion follows from Proposition 5.3. Now suppose $\nu\geq 3$, and the assertion holds for all $\nu-1$. If $N^{G}(\mathcal{H})\neq\{1\}$, take a subgroup $N$ of order $p$ in $Z(G)\cap N^{G}(\mathcal{H})$. Note that $Z(G)\cap N^{G}(\mathcal{H})$ is non-trivial by Lemma 4.3. Then it suffices to prove the assertion for the $G/N$-lattice $J_{(G/N)/\mathcal{H}^{N}}$. Since $(G/N)=p^{\nu-1}$, the assertion follows from the induction hypothesis. Hence, we may further assume $N^{G}(\mathcal{H})=\{1\}$. If $M(\mathcal{H})=p$, then $G$ is elementary $p$-abelian since $N^{G}(\mathcal{H})=\{1\}$. Therefore, the $G$-lattice $J_{G/\mathcal{H}}$ is not quasi-invertible by Proposition 5.3. Therefore, we can impose $M(\mathcal{H})\geq p^{2}$ in the sequel. By the induction hypothesis, it suffices to prove that there is a maximal subgroup $P$ of $G$ that satisfies $\mathcal{H}_{P}^{\operatorname{red}}\geq 2$.
@@ 6.1. Quasi-permutation lattices
where the horizontal sequences are exact. Note that the central and the rightmost vertical maps are injective, and the cokernel of the central vertical map is torsion-free. Moreover, since the cokernel of $\Phi_{m,\nu,\ast}$ is torsion, we have
@@ 7.3. Third step: The case admitting normal factors
Case 2-c. $\Sigma=\sigma$. Let $P$ be a maximal subgroup of $G$ containing $N_{G}(H_{1})$. It suffices to prove that $\mathcal{H}_{P}^{\operatorname{red}}$ is not a subset of $\{gH_{1}g^{-1}\mid g\in G\}$. This implies $\mathcal{H}_{P}^{\operatorname{red}}\geq 3$ since $H_{1}$ is not normal in $G$. Assume not, that is, $\mathcal{H}_{P}^{\operatorname{red}}$ is contained in $\{gH_{1}g^{-1}\mid g\in G\}$. Since $H_{0}$ is normal of index $\Sigma=\sigma=\mu(\mathcal{H})$, it is contained in $N_{G}(H_{1})$ by Lemma 5.9. In particular, $H_{0}$ is an element of $\mathcal{H}_{P}^{\operatorname{red}}$, which contradicts the assumption. Hence the proof is complete. $\blacksquare$
@@ 7.4. Final step: General case
      $G/N^{G}(\mathcal{H})\cong D_{2^{\nu}}$, $\mathcal{H}=\{H,H^{\prime}\}$, $H/N^{G}(\mathcal{H})\cong\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle$ and $H^{\prime}/{N^{G}(\mathcal{H})}\cong\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle$ for some $\nu\in\mathbb{Z}_{>0}$ and $m,m^{\prime}\in\mathbb{Z}$.
@@
(ii) $\Rightarrow$ (iii): Since $G$ is nilpotent, we have $G=\prod_{p\mid\#G}G_{p}$, where $G_{p}$ is a $p$-Sylow subgroup of $G$. Hence, applying Lemma 8.3 for all prime divisors of #G\#G, we obtain a reduced set $\mathcal{H}^{\prime\prime}$ of subgroups in $G$ that satisfies
  • $\mathcal{H}^{\prime\prime}<\mathcal{H}$;
  • $(\mathcal{H}^{\prime\prime}_{G_{p}})^{\operatorname{set}}=\mathcal{H}_{G_{p}}^{\operatorname{red}}$ for all $p$; and
  • $J_{G/\mathcal{H}^{\prime\prime}}]=[J_{G/\mathcal{H}}]$.
On the other hand, since $J_{G/\mathcal{H}}$ is quasi-invertible as a $G_{p}$-lattice, Theorem 5.10 implies that $(\mathcal{H}^{\prime\prime}_{G_{p}})^{\operatorname{set}}$ consists of a single normal subgroup $N_{p}$ of $G_{p}$ with $G_{p}/N_{p}$ cyclic. Since $\prod_{2<p\mid\#G}N_{p}$ is normal in $G$, we may assume that $G_{p}$ is cyclic and $N_{p}=\{1\}$ for any odd prime divisor $p$ of #G\#G. In particular, we have
$G\cong C_{m}\times G_{2}$
for some odd positive integer. Moreover, #H\#H is a power of $2$ for any $H\in\mathcal{H}^{\prime\prime}$. Now, set
$\mathcal{H}^{\prime}:=(\mathcal{H}^{\prime\prime})^{\operatorname{srd}}.$
Then $(\mathcal{H}^{\prime}_{G_{2}})^{\operatorname{set}}$ is also strongly reduced.
Case 1. $\mathcal{H}^{\prime}=1$.
In this case, Theorem 7.16 implies that $(\mathcal{H}^{\prime}_{G_{2}})^{\operatorname{set}}$ satisfies (iii-I), that is, $(\mathcal{H}^{\prime}_{G_{2}})^{\operatorname{set}}$ consists of a normal subgroup $N$ of $G_{2}$ with $G_{2}/N$ cyclic. Moreover, we have $N\triangleleft G$ and $G/N$ is cyclic. Therefore, $\mathcal{H}^{\prime}$ satisfies (iii-$\alpha$).
Case 2. $\mathcal{H}^{\prime}\geq 2$.
In this case, Theorem 7.16 asserts that (iii-II) is valid for $(\mathcal{H}^{\prime}_{G_{2}})^{\operatorname{set}}$, that is,
  • $G_{2}/N^{G_{2}}(\mathcal{H})\cong D_{2^{\nu}}$ for some $\nu\in\mathbb{Z}_{>0}$;
  • $\mathcal{H}=\{H,H^{\prime}\}$;
  • $H/N^{G}(\mathcal{H})\cong\langle\sigma_{2^{\nu}}^{2m}\tau_{2^{\nu}}\rangle$ and $H^{\prime}/{N^{G}(\mathcal{H})}\cong\langle\sigma_{2^{\nu}}^{2m^{\prime}+1}\tau_{2^{\nu}}\rangle$ for some $m,m^{\prime}\in\mathbb{Z}$.
Since $N^{G_{2}}(\mathcal{H}^{\prime})$ is normal in $G$, we have $N^{G}(\mathcal{H})=N^{G_{2}}(\mathcal{H})$ and $G/N^{G}(\mathcal{H})\cong C_{m}\times D_{2^{\nu}}$. Consequently, $\mathcal{H}^{\prime}$ satisfies (iii-$\beta$).
@@
(ii) $\Rightarrow$ (iii): Combining the assumption with Proposition 3.1, we obtain that the $G$-lattice $J_{G/\mathcal{H}}$ is quasi-invertible. Hence, there exists a reduced set of subgroups in $G$ with $\mathcal{H}^{\prime}<\mathcal{H}$ and $J_{G/\mathcal{H}^{\prime}}]=[J_{G/\mathcal{H}}]$ that satisfies (iii-$\alpha$) or (iii-$\beta$) in Theorem 8.4. If (iii-$\alpha$) holds, denote by $\mathbf{K}^{\prime}=K^{\prime}$ the intermediate field of $\widetilde{K}/k$ corresponding to $N$. Then, $K^{\prime}$ is a subalgebra of $\mathbf{K}$ satisfying (iii-$\alpha$) in Theorem 1.5. On the other hand, if (iii-$\beta$) is true, write $\mathcal{H}=\{H,H^{\prime}\}$, and denote by $K^{\prime}_{1}$ and $K^{\prime}_{2}$ the intermediate fields of $L/k$ corresponding to $K^{\prime}_{1}$ and $K^{\prime}_{2}$, respectively. Let $\mathbf{K}^{\prime}:=K^{\prime}_{1}\times K^{\prime}_{2}$, then it is a subalgebra of $\mathbf{K}$ satisfying (iii-b) in Theorem 1.5.
@@
$^{1}(k,T):=\operatorname{Ker}\left(H^{1}(k,T)\xrightarrow{(\operatorname{Res}_{k_{v}/k})_{v}}\prod_{v}H^{1}(k_{v},T)\right).$
@@
Here, $\operatorname{Pic}(\overline{X})^{\vee}$ and $A_{k}(T)$ are defined as follows: